a 10% solution of fertilizer is to be mixed with a 70% solution of fertilizer in order to get 120 gallons of a 50% solution. How many gallons of the 10% and 70% solution should be mixed?
Answer by TimothyLamb(4379)

---
x = gallons of 10% mix
y = gallons of 70% mix
---
10x + 70y = 50*120
x + y = 120
---
put the system of linear equations into standard form:
---
10x + 70y = 6000
x + y = 120
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = gallons of 10% mix = 40
y = gallons of 70% mix = 80
---
system matrix
10 70
1 1
---
inverse of system matrix
-0.016666667 1.1666667
0.016666667 -0.16666667
---
determinant of system matrix = -60
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php